OpenStudy (anonymous):

Factor the trinomial and write each factor as a polynomial in descending order: 12x^2+20x-25

myininaya (myininaya):

12(-25)=6(-50)=30(-10) 30-10=20 so we have 12x^2+30x-10x-25 =6x(2x+5)-5(2x+5) =(2x+5)(6x-5)

OpenStudy (aroub):

(16x+5)(6x-5)

OpenStudy (aroub):

wow myininaya i loved the way how do u find it! :D EASY!

myininaya (myininaya):

lol :)